A sample of an arsenic salt weighing 100mg was dissolved and the arsenic converted to arsenite, which required 10.98 mL of 0.01234 M permanganate solution for complete reaction. What is the percentage of arsenic in the salt? The atomic weight of arsenic is 74.92 g mol -1.
Answer #1

Number of moles of Permanganate solution = 10.98/1000 * 0.01234 = 0.00013549 moles

Arsenic will go from oxidation state of +3 to +5, hence number of moles of arsenic will be

0.00013549 moles * 5/2 = 0.000338733 moles

Mass of arsenic = 0.000338733 moles * 74.92 gm/mol = 0.025377 gms

Percentage of arsenic in salt = 0.025377/0.1 * 100 = 25.37%